OpenStudy (brittanyrolfe):

For the ideal gas law, what is the R constant in the formula: PV= nRT? When P is in atm, R is in liters, n is number of moles and T is in Kelvins.

OpenStudy (brittanyrolfe):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

So the Ideal gas law corresponds to the ideal gas equation, which you laid out very nicely here. So R is representative of the ideal gas constant. R is equal to \(0.08206 L \frac{ L \times atm}{mol \times K}\) Where L is liters (volume), K is Kelvin (temperature), and atm is atmospheric pressure.